The Solar and Heliospheric Observatory (SOHO) spac
Answer #1

r^3 = GMT^2/4pi^2

r^3 = 6.67*10^-11*5.974*10^24*(365*24*60*60)^2/4*3.14^2

r^3 = 100048.12*10^24

r = 21.579*10^8 m

R = 2.16*10^9-earth raius